Find the probabilities using the normal distribution



1. P (0 <z <1.96)


2. P (z> -1.43)


3. P (1.12 <z <1.43)


4. P (-1.23 <z <0)

1
Expert's answer
2022-03-29T06:05:41-0400

1) P(0<Z<1.96)=P(Z<1.96)-P(Z<0)=0.4750.

2) P(Z>-1.43)=P(Z<1.43)=0.9236.

3) P(1.12<Z<1.43)=P(Z<1.43)-P(Z<1.12)=0.0550.

4) P(-1.23<Z<0)=P(0<Z<1.23)=P(Z<1.23)-P(Z<0)=0.3907.